OpenStudy (anonymous):

The area of a rectangle can be found by multiplying the length by the width a rectangular postage stamp has a length of nine-tenths inch and a width of five-sixths inch what is the area of the stamp in square inches?

OpenStudy (anonymous):

area=9/10 times 5/6 area = 45/60 area=0.75in^2

OpenStudy (imstuck):

or\[\frac{ 9 }{ 10 }*\frac{ 5 }{ 6 }\]the 5 would cancel with the 10, making the 10 a 2, and 3 goes into both 9 and 6, leaving\[\frac{ 3 }{ 4 }"^{2}\]Coouldn't get that to say "inches squared" so I did that.

OpenStudy (anonymous):

thank you